OpenStudy (anonymous):

Hey, you guys bored? :D\[\int_0^af\left(x\right)\,dx=\int_{-a}^afrac{f\left(x\right)}{1+e^x}\,dx\]Show that the above is true under the conditions \(f:\mathbb{R}\rightarrow\mathbb{R},f\,\text{is continuous},a\in\mathbb{R},a>0\). Good luck!

OpenStudy (anonymous):

\int_0^af\left(x\right)\,dx=\int_{-a}^a\frac{f\left(x\right)}{1+e^x}\,dx

OpenStudy (turingtest):

\[\int_0^af\left(x\right)\,dx=\int_{-a}^a\frac{f\left(x\right)}{1+e^x}\,dx\]

OpenStudy (anonymous):

Yes!

OpenStudy (turingtest):

interesting-looking :)

OpenStudy (anonymous):

Oh, sorry, forgot to specify \(f\,\text{is even}\).

OpenStudy (turingtest):

I was about to say there was some even business happening here that will help I'm sure

OpenStudy (turingtest):

1/(1+e^x) is not even

OpenStudy (anonymous):

How is \(\displaystyle{\frac{f\left(x\right)}{1+e^x}}\) even?

OpenStudy (kinggeorge):

You're right. It's odd around \(x=0\), \(y=.5\) My bad. Ignore my last comment.

OpenStudy (anonymous):

Rewriting the entire question under one post, just cause. Under the conditions:\[f:\mathbb{R}\rightarrow\mathbb{R},f\,\text{is continuous},f\left(x\right)=f\left(-x\right),a\in\mathbb{R},a>0\]Prove that:\[\int_0^af\left(x\right)\,dx=\int_{-a}^a\frac{f\left(x\right)}{1+e^x}\,dx\]

OpenStudy (turingtest):

I am trying a trig sub and interesting things are happening...

OpenStudy (zarkon):

hint: \[\int\limits_{-a}^{a}\frac{f(x)}{1+e^{x}}dx=\int\limits_{-a}^{0}\frac{f(x)}{1+e^{x}}dx+\int\limits_{0}^{a}\frac{f(x)}{1+e^{x}}dx\]

OpenStudy (zarkon):

u=-x

OpenStudy (anonymous):

\[\int_a^bf\left(x\right)\,dx=\int_a^bf\left(a+b-x\right)\,dx\]This is an important identity for the solution, apparently.

OpenStudy (turingtest):

I want to use zarkons tip, I bet it works I feel I am making progress

OpenStudy (anonymous):

I don't suppose anyone knows what this identity is? I'd like to read up its derivation.

OpenStudy (zarkon):

\[u=-x\] \[du=-dx\] \[\int\limits_{-a}^{0}\frac{f(x)}{1+e^x}dx\] \[=-\int\limits_{a}^{0}\frac{f(-u)}{1+e^{-u}}du\] \[=\int\limits_{0}^{a}\frac{f(u)}{1+e^{-u}}du\] \[=\int\limits_{0}^{a}\frac{f(x)}{1+e^{-x}}dx\] \[\int\limits_{-a}^{a}\frac{f(x)}{1+e^{x}}dx=\int\limits_{-a}^{0}\frac{f(x)}{1+e^{x}}dx+\int\limits_{0}^{a}\frac{f(x)}{1+e^{x}}dx\] \[=\int\limits_{0}^{a}\frac{f(x)}{1+e^{-x}}dx+\int\limits_{0}^{a}\frac{f(x)}{1+e^{x}}dx\]

OpenStudy (zarkon):

\[=\int\limits_{0}^{a}\frac{f(x)}{1+e^{-x}}dx+\int\limits_{0}^{a}\frac{f(x)}{1+e^{x}}dx\] \[=\int\limits_{0}^{a}\left[\frac{f(x)}{1+e^{-x}}+\frac{f(x)}{1+e^{x}}\right]dx\] \[=\int\limits_{0}^{a}\left[\frac{f(x)(1+e^x)+f(x)(1+e^{-x})}{(1+e^{-x})(1+e^x)}\right]dx\] \[=\int\limits_{0}^{a}\left[\frac{f(x)(2+e^{x}+e^{-x})}{2+e^x+e^{-x}}\right]dx\] \[=\int\limits_{0}^{a}f(x)dx\]

OpenStudy (turingtest):

sorcery I say! Zarkon rocks :)